Number of words of 4 letters that can be formed with the letter of the word IITJEE is...............


 

Grade:12

14 Answers

Maruthi Sirisilla
10 years ago

the answer is 468

Anshuman Chandel
10 years ago
102
shetti swamy patel
10 years ago

answer is 256

shubham sharma
10 years ago

Ans. Is 14!!!

» four distinct letters= (4C4)

»Of form (aabb)= (2C1).(4C2)

»Of form aabc= (2C2)

So 0n adding it comes 14.

 

revathy t.nair
10 years ago

is that 6*6*6*6=1296? is that the answer?

Sandeep Trivedi
10 years ago

72

jitender lakhanpal
10 years ago

there will be following cases

1) words form by letters I,T,J,E

we will use permutation

4! =24

2) words form by only I,I,E,E

there would be following arrangement

IIEE   another word by flipping the letters I,E

IEEI another word by flipping the letters

EIEI  another word by flipping the letters

it makes total of 6

3) words formed by 2 I''S T,J,E

first select 2 places out of four so we use combination

3C2  and then arrange 3 letters in 2 places

so finally we get

4C2*6 = 36

4) words formed by 2 E''S  AND T,J,E

so again 36

so add all cases you will get the answer

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)

Krrish .
10 years ago
184
ashwin26
10 years ago
IITJEE Can be written as " II ", "T" , "J" , "EE" here there are two sets of repeated combinations(I,J) .now the four letter words can be categorised into (i)two similar,another two similar 2C1*(4!/(2!*2!)) ways. (ii)two similar , 1 different , another different 2C1*(3C2)*(4!/2!) ways (iii)All different 4! ways so the total number of 4 letter words= 12+72+24=108
GOPI
10 years ago
90
KARTHIK
10 years ago
there will be following cases 1) words form by letters I,T,J,E we will use permutation 4! =24 2) words form by only I,I,E,E there would be following arrangement IIEE another word by flipping the letters I,E IEEI another word by flipping the letters EIEI another word by flipping the letters it makes total of 6 3) words formed by 2 I``S T,J,E first select 2 places out of four so we use combination 3C2 and then arrange 3 letters in 2 places so finally we get 4C2*6 = 36 4) words formed by 2 E``S AND T,J,E so again 36 so add all cases you will get the answer
venkatesh vadde
10 years ago
24
venkatesh vadde
10 years ago
no restriction means digits can be repeated if first place is filled by I then possible arrangements are 5!/2! if first place is filled by T then possible arrangements are 5!/(2!*2!) if first place is filled by J then possible arrangements are 5!/(2!*2!) if first place is filled by E then possible arrangements are 5!/2! total no of arrangements are 60+30+30+60=180
goutham
8 years ago
102